[figure: 9.7 cm]

Switching of magnetostriction in rare earth Ising metamagnet induced by small sign-variable increments of a magnetic field.

I. B. Krynetskii^{a},V. M. Matveev^{b},V. V. Matveev^{b} aPhys. Depart. Moscow State University, 119992, Moscow, Russia bZelenograd Research Institute of Physical Problems, 103460, Moscow, Russia

Experimental and theoretical studies of magnetostriction for Ising metamagnet DyAlO_{3} were performed in the temperature range of 1.5 − 4.2 K and magnetic fields up to 4 T. The magnetostriction curves measured during ”field-up” and ”field-down” magnetisation processes was found to be near-symmetric about the H = 0 axis and form a peculiar magnetostriction hysteresis loop. Besides, at any value of the magnetic field (ranging from very small to high fields) small sign-variable increments of the magnetic field result in reversible switching of magnetostriction between the ”field-up” and ”field-down” curves. Such a behaviour is similar to that of the magnetisation curves for ”hard” type-II superconductors with strong pinning. Possible mechanisms of these anomalies are proposed. A layered magnetic structure of the dipole Ising antiferromagnet DyAlO_{3} and the existence of a great number of metastable magnetic phases which are near-equal in energy and break up the crystal into ”quasi-mesoscopic” regions play a key role in the observed anomalies. The work was supported by the RFBR, Grant No 02-02-16704.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : I. B. Krynetskii

Address for correspondence : Physics Department, Lomonosov Moscow State University, 119992, Moscow, Russia

Email address : krynets@plms.phys.msu.su